OpenStudy (anonymous):

passes through(4,2) and (3,-2)?show working

OpenStudy (anonymous):

let points be (a1,b1)and (a2,b2)........then substitute in the formulae......y-a2=(b2-b1/a2-a1)(x-a2)............

OpenStudy (anonymous):

is the question to find the gradient of the line that passes through those points?

OpenStudy (anonymous):

if it is, then it's: \[\frac{ 2-(-2) }{ 4-3 }\] which simplifies to \[\frac{ 4 }{ 1 }\] which equals 4.

OpenStudy (anonymous):

basically, to find the gradient of a line, you use the formula: \[\frac{ y _{1}-y _{2} }{ x _{1} -x _{2}}\]